OpenStudy (anonymous):

Need Help with these 4 questions Please. 1. Write the equation of a line in point slope form that has a slope of 43− and passes through the point (4, 5). 2. Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4). 3. Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5) 4. Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

1. y-5=43(x-4) 2. y=4 3. don't know 4. y=4x+53

OpenStudy (igreen):

For questions like #3, I just use https://www.desmos.com/calculator and experiment.

OpenStudy (igreen):

So for question #3, your equation is y = 3x + 32.5.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Stephany_Velez